hello,
can anyone help me solve these linear problems ? I really need help & im completely clueless about how to do this . Help ?
1. Wade purchased three videos and one music CD. The CD cost Wade $12.99. If he paid the same amount for each video and spent total of $42.96, how much did each video cost ? ( show work please ? )
2. At the market, Xiang bought a bunch of bananas for$0.35 per pound and frozen pizza for $4.99. The total for Xiang's purchase was $0.64 without tax. How many pounds of bananas did Xiang buy? ( show work please ? )
3. Yamir went to the store to buy gardening supplies. A bag of dirt was $3.99 and tulips cost 75 cents per bulb. He bought one bag of dirt and some tulip bulbs and spent a total of $12.24 without tax. How many bulbs did Yamir buy?
( show work please ? )
4. Zoe is comparing two local yoga programs. Yoga-Weigh charges $90 dollars a month and a registration fee of $35. Essence of Yoga charges $80 per month with $75 registration fee. After how many months will the two schools charge the same amount ? ( show work please ? )

Answer by umnikmo(75)

1. price of a video is x, so
12.99+3x=42.96, from here x=(42.96-12.99)/3=9.99
2.number of pounds of bananas is N, so
0.35*N+4.99=(insert correct value, 0.64 does not look right), then solve for N
3.number of bulbs - z
3.99+.75z=12.24
from here z=(12.24-3.99)/0.75=11
4.C1- cost of Yoga-Weigh for M months
C2 - cost of Essence of Yoga for M months
C1=35+90M
C2=75+80M
these two costs will be the same, C1=C2, so
35+90M=75+80M
10M=40
M=4
ans: after 4 months two schools will charge same ammount